XYZ funds purchased a bank bill (face value=1 million) on 10th June 2020. The bill will mature in 60 days and the yield is 6.95%. After holding the bill for 40 days, the funds sold the bill at a yield of 6.5%.


a) Calculate the profit through buy and sell of this bank bill. (4.5 marks)

b) Work out the simple annual interest rate and effective annual interest rate earned by XYZ.(4 marks)


1
Expert's answer
2021-05-11T14:16:00-0400

a).

First, purchase price will be calculated.

Face value = 1,000,000

Yield rate = 6.95%

Time = 60 days

"Face\space value=1,000,000"

"Discounted\space amount=1,000,000\times9.65\%\times\frac{60}{365}"

"=11,424.66"

"Purchase\space price=Face\space value-Discounted\space amount"

"=1,000,000-11,424.66"

"=988,575.37"


"Time=20\space days(60-40)"

"Yield\space on\space sale=6.5\%"

"Sale\space price=1,000,000-(1,000,000\times 6.5\%\times\frac{20}{365})"

"=1,000,000-3,561.64"

"=996,438.36"


"Profit\space earned=Sale\space price-Purchase\space price"

"=996,438.36-988,575.34"

"=7,863.02"


Hence profit is "7,863.02"


(b)

Computation of interest:-

"\frac{Simple\space interest}{Profit}=Amount\space invested\times Rate\times Time"

"7,863.02=988,575.34\times Rate \times\frac{40}{365}"

"7,863.02=108,337.02\times Rate"

"Rate=\frac{7,863.02}{108,337.02}"

"=0.07258\space or\space 7.158\%"

"Effective\space annual\space interest\space rate(EAR)"


"=(1+\frac{Rate}{compound\space period})^{compounding\space period}-1"


"=(1+\frac{0.07258}{365})^{365}-1"


"=(1+0.000199)^{365}-1"



"=1.07533-1"


"=0.07533\space or\space 7.533\%"


Simple interest is 7.158% and EAR is 7.533%